OpenStudy (camerondoherty):

Which shows 232 - 172 being evaluated using the difference of perfect squares method? 232 - 172 = (529 + 289)(529 - 289) = 196,320 232 - 172 = 529 - 289 = 240 232 - 172 = (23 + 17)(23 - 17) = (40)(6) = 240 232 - 172 = (23 - 17)2 = (6)2 = 36

OpenStudy (anonymous):

its B...=529-289=(23)^2-(17)^2

OpenStudy (anonymous):

to me, none of them. :( it's quite simple, no matter what method is, the final solution must be 60 , but non of the options is 60 :( I fail this question

OpenStudy (camerondoherty):

*confused*

OpenStudy (camerondoherty):

@marissalovescats

OpenStudy (dan815):

\[23^2-17^2=(23+17)(23-17) = 40*6=240\]

OpenStudy (anonymous):

what the.... hahahaa 23^2 not 232